# Compression of quantum shallow-circuit states

### v1

@@ III.2 Proof of Theorem 1
As stated in Protocol 1 (step 6), the classical memory is to store $\hat{U}$, an element of the covering $\mathsf{U}_{{\rm sc}}(\epsilon_{0})$. By Lemma 9 and substituting in $\epsilon_{0}=n\cdot d\cdot N^{-\frac{1}{2}(1-\Delta)}$ and $G=nd/2$, the size of the classical memory is bounded as
$M_{\rm c}=\log_{2}|\mathsf{U}_{\rm sc}(\epsilon_{0})|\leq 8(1-\Delta)nd\left(\log_{2}N+O\left(\log_{2}n\right)\right).$ (88)
The quantum memory is used to store the truncated multi-mode states. Again, since the truncation for each mode costs no more than $\log_{2}(e\alpha_{0})^{2}$ qubits (with $\alpha_{0}=896ndN^{\Delta/2}$) by Lemma 8 and there are no more than $(n/d+4)2^{8d}$ modes, the size of the quantum memory is bounded as
$M_{\rm q}\leq(n/d+4)2^{8d}\cdot\left(\Delta\log_{2}N+O(\log_{2}n)\right).$ (89)
V.5 Optimality of the memory scaling.
Here we show that the memory consumption of Protocol 1 is the minimum at least in scaling.
@@
V.6 Necessity of quantum memory.
In Subsection V.3, we have seen that the compression does not require a fully quantum memory. Even more, under certain conditions, the hybrid memory can be made almost classical. It is thus tempting to ask whether quantum memory is necessary at all, i.e., whether a fully classical memory can fulfill the goal of faithfully compression. Here we give a negative answer by showing that any protocol using only classical memory cannot achieve faithfulness. Note that this fact was first shown in Ref. [33] for qudits; see Proposition 1. Here we show the same fact for shallow-circuit states by making some modifications to the original proof.
@@
$\max_{j=1,2}\left\{d_{\operatorname{Tr}}(\mathcal{V}_{N}(\psi_{j}),G_{j}),d_{\operatorname{Tr}}(\psi_{j},\mathcal{V}_{N}^{\ast}(G_{j})\right\}=:\epsilon_{{\rm Q-LAN},N}\qquad\lim_{N\to\infty}\epsilon_{{\rm Q-LAN},N}=0.$ (113)
@@
where $\mathcal{V}_{N}$ and $\mathcal{V}_{N}^{\ast}$ are the Q-LAN and the inverse Q-LAN, respectively (cf. Section IV).
@@
V.7 Computational complexity of the compression.
We conclude this section with a brief discussion on the computational efficiency of Protocol 1. To execute the protocol on a quantum computer, one needs to execute the tomography subroutine, the Q-LAN and inverse Q-LAN subroutines, and the truncation subroutine. The complexity analysis can be done separately for each subroutine.
For the Q-LAN (and its inverse), we have to execute the channels (63) and (64). Since we are dealing with pure states only, it is enough to implement the transformation from the uncoupled basis to the symmetric basis. For instance, this can be realized via the Schur transform (which is a bit of overkill since we are interested only in the symmetric subspace). According to Refs. [54, 55, 56], the Schur transform can be done with polynomial gate complexity in both $N$ and $\log_{2}D$ (with $D:=2^{n}$ being the single-copy dimension), and thus our Q-LAN subroutines are computationally efficient. The truncation subroutine requires repeating $\mathsf{poly}(n)$ times the same simple truncation on a single mode (register), which can be done efficiently with the assistance of ancilla. However, the tomography subroutine of Protocol 1 is not computationally efficient. The main reason is that it requires hypothesis selection among exponentially many candidates; see Subsection V.2.
Based on the above discussion, Protocol 1 requires exponential gate complexity in $n$, with the bottleneck being the hypothesis selection at the tomography step. This might be improved in future works by lifting the reliance on the $\epsilon$-covering in the tomography subroutine.
Acknowledgement. The author acknowledges Penghui Yao and Fang Song for helpful discussions on pseudorandom states. This work is supported by the National Natural Science Foundation of China via Excellent Young Scientists Fund (Hong Kong and Macau) Project 12322516, Guangdong Provincial Quantum Science Strategic Initiative (GDZX2303007 and GDZX2200001), Guangdong Basic and Applied Basic Research Foundation (Project No. 2022A1515010340), and the Hong Kong Research Grant Council (RGC) through the Early Career Scheme (ECS) grant 27310822 and the General Research Fund (GRF) grant 17303923.

### v3

@@ B.2 Proof of Theorem 4
As stated in Protocol 2 (step 6), the classical memory is to store $\hat{U}$, an element of the covering $\mathsf{U}_{{\rm sc}}(\epsilon_{0})$. By Lemma 9 and substituting in $\epsilon_{0}=n\cdot d\cdot N^{-\frac{1}{2}(1-2\Delta/3)}$ and $G=nd/2$, the size of the classical memory is bounded as
$M_{\rm c}=\log_{2}|\mathsf{U}_{\rm sc}(\epsilon_{0})|\leq 8\left(1-\frac{2\Delta}{3}\right)nd\left(\log_{2}N+O\left(\log_{2}n\right)\right).$ (88)
The quantum memory is used to store the truncated multi-mode states. Again, since the truncation for each mode costs no more than $\log_{2}(e\alpha_{0})^{2}$ qubits (with $\alpha_{0}=896ndN^{\Delta/3}$) by Lemma 8 and there are no more than $(n/d+4)2^{8d}$ modes, the size of the quantum memory is bounded as
$M_{\rm q}\leq\frac{(n/d+4)2^{8d+1}}{3}\left(\Delta\log_{2}N+O(\log_{2}n)\right).$ (89)
D.5 Optimality of the memory scaling.
Here we show that the memory consumption of Protocol 2 is the minimum at least in scaling.
@@
D.6 Necessity of quantum memory.
In Appendix D.3, we have seen that the compression does not require a fully quantum memory. Even more, under certain conditions, the hybrid memory can be made almost classical. It is thus tempting to ask whether quantum memory is necessary at all, i.e., whether a fully classical memory can fulfill the goal of faithfully compression. Here we give a negative answer by showing that any protocol using only classical memory cannot achieve faithfulness. Note that this fact was first shown in Ref. [33] for qudits; see Proposition 1. Here we show the same fact for shallow-circuit states by making some modifications to the original proof.
@@
$\max_{j=1,2}\left\{d_{\operatorname{Tr}}\left(\mathcal{V}_{N}(\psi_{j}^{\otimes N}),G_{j}\right),d_{\operatorname{Tr}}\left(\psi_{j}^{\otimes N},\mathcal{V}_{N}^{\ast}(G_{j})\right)\right\}=:\epsilon_{{\rm Q-LAN},N}\qquad\lim_{N\to\infty}\epsilon_{{\rm Q-LAN},N}=0.$ (113)
@@
where $\mathcal{V}_{N}$ and $\mathcal{V}_{N}^{\ast}$ are the Q-LAN and the inverse Q-LAN, respectively (cf. Appendix C).
@@
D.7 Time complexity and gate complexity of the compression.
As shown in the description of Protocol 2, the compression consists of several steps. Their execution times are analyzed in the following:
  • Tomography. The tomography subroutine [9] we plan to use has a gate complexity that is at least exponential in $G$, the number of gates in the generating circuit of the state. Instead of directly executing the tomography subroutine [9], we make it more computationally efficient by using the structure of shallow circuits. Since the brickwork shallow circuit can be reduced into two layers (cf. Section B), the tomography can be done in just two consecutive steps. First, gates in the first layer are determined (up to local unitaries) with a small error. Since all gates in the same layer commute with each other, they can be determined in parallel, and thus the time complexity is equal to the complexity for a single gate (with an error that vanishes at most as $\mathsf{poly}(n)$). Next, the same procedure is applied to gates in the second layer. Since each gate is of a constant size and there are $O(n)$ gates in each layer, the whole tomography can be done efficiently.
  • The Q-LAN and the inverse Q-LAN. Essentially, one needs to implement any transformation such that Eq. (61) (and its pseudo-inverse) holds, which is the transformation from a certain subset of computational states to the corresponding symmetric states.
    Here we present a concrete implementation of this step consisting of two subroutines. The first is to apply the same unitary transformation $F$ on each $n$-qubit system that reorders the computational bases by the number of ones they contain: $F$ maps a computational basis state $|x\rangle$ into $|f_{x}\rangle$, where $f_{x}$ is a binary string such that $f_{x}\geq f_{y}$ if there are more ones in $x$ than in $y$. This $F$ is a crucial step of the Schumacher compression and can be achieved by the construction in Ref. [68], which has time complexity $O(n^{3})$ with $O(\sqrt{n})$ ancillary qubits. In total, $N$ uses of $F$ in parallel, and thus this subroutine is both time-efficient and computationally efficient. For the inverse Q-LAN in decoding, $F^{\dagger}$ can be implemented (efficiently) with the same complexity. The purpose of doing so is to inversely order the basis states by the number of ones for the convenience of the truncation (see the next step).
    The second subroutine is a basis transformation into the symmetric basis. It is sufficient (but not necessary) to implement the Schur transform (and its inverse) on the $N$ systems. According to Refs. [69, 70, 43], the Schur transform can be done with polynomial gate complexity in both $N$ and $\log_{2}D$ (with $D:=2^{n}$ being the single-copy dimension), and thus our Q-LAN subroutines are computationally efficient and, as an immediate consequence, time efficient. The algorithm in Ref. [43] outputs the representation register in the Gelfand-Tsetlin basis. More explicitly, it takes states in the computational basis as inputs and outputs states in the Schur basis, which has three registers: the index register (labeled by Young diagrams), the multiplicity register, and the representation register. Since only symmetric states are concerned here, the first two registers are trivial, and the third is spanned by the Gelfand-Tsetlin basis. For symmetric states of $\mathcal{H}_{D=2^{n}}^{\otimes N}$ with type $(N_{1},\dots,N_{2^{n}})$, the nonzero subregister of the Gelfand-Tsetlin state reads $|N,N-N_{1},N-N_{1}-N_{2},\dots,N_{2^{n}}\rangle$. Here $N_{j}$ corresponds to the number of $|j\rangle$.
  • Photon-number truncation. We start from the Gelfand-Tsetlin state which reads $|N,N-N_{1},N-N_{1}-N_{2},\dots,N_{2^{n}}\rangle$ with $2^{n}$ sub-registers. Crucially, note that now the computational basis has been reordered by the unitary $F$ at the previous step (cf. the first subroutine), and now $N_{j}$ corresponds to the original binary string with the $j$-th lowest number of ones. By Eq. (58), we can focus only on the first $O(2^{d}n)$ sub-registers with at most $O(d)$ ones adjacent to each other, i.e., those corresponding to binary strings in $\mathsf{B}_{j}$ (56). One can simply discard the other $O(2^{n})$ sub-registers. The remaining registers (consisting of $O(2^{n}n)$ sub-registers, each of dimension $N$) $|N,N-N_{1},N-N_{1}-N_{2},\dots\rangle$ can be transformed into a type register $|N_{1},N_{2},\dots\rangle$ by consecutive subtraction, which can be done efficiently.
    Next, we need to do a number truncation on each sub-register ($N$-dimensional), as required in Lemma 8. Since the truncation operations on different modes are independent and can be implemented in parallel. It is sufficient to analyze the time it takes to implement one truncation.
    Suppose one is given a register (consisting of qubits) in a pure state. In our setup, to realize a truncation above (or below, in a similar way) a threshold value $N_{0}$, it is enough to simply discard all but the last $\lceil\log_{2}(N_{0}-1)\rceil$ qubits. Indeed, as the state can be represented as
    $|\psi\rangle=|0\cdots 0\rangle\otimes\sum_{b\leq N_{0}}\psi_{b}|b\rangle+\sqrt{\epsilon_{0}}|\psi^{\prime}\rangle\in\mathcal{H}_{A}\otimes\mathcal{H}_{B},$ (127)
    for some normalized vector $|\psi^{\prime}\rangle$ orthogonal to $\{|0\cdots 0\rangle\otimes|b\rangle\}$, where $\mathcal{H}_{B}$ corresponds to the last $\lceil\log_{2}(N_{0}-1)\rceil$ qubits and $\mathcal{H}_{A}$ corresponds to the other qubits, $b$ is the basis index in the binary representation and $\psi_{b}$ denotes the corresponding amplitude with $\sum_{b\leq N_{0}}|\psi_{b}|^{2}=1-\epsilon_{0}$. Discarding the first sub-register yields $\rho=\sum_{b,b^{\prime}\leq N_{0}}\psi_{b}\psi^{\ast}_{b^{\prime}}|b\rangle\langle b^{\prime}|+\operatorname{Tr}_{A}\psi^{\prime}$, which is a state on $\lceil\log_{2}(N_{0}-1)\rceil$ qubits, and appending zero qubits to it yields a state $|0\cdots 0\rangle\langle 0\cdots 0|\otimes\rho$ with at least $(1-\epsilon_{0})^{2}$ fidelity with $\psi$. Therefore, the truncation would have a negligible error if $\epsilon_{0}$ is small.
    Overall, the truncation can be done time efficiently. Whether it is also gate-efficient depends on the complexity of discarding/resetting a sub-register. As one has to discard $O(2^{n})$ sub-registers (in parallel, so the time complexity is low), the truncation step is gate-efficient if the complexity cost of discarding these registers is zero; otherwise, this step is not gate-efficient.
  • Amplification. For amplification, we need to apply the same single-mode amplifier to each of the $O(n)$ modes. Again, we may simply analyze the efficiency of a single-mode amplifier. The amplification can be realized by the original proposal in Ref. [60], where the mode is jointly squeezed together with a vacuum mode by a two-mode squeezing operation. The degree of squeezing is related to the degree of amplification required and is close to one in our case (i.e., the two-mode squeezing operation is close to identity). The two-mode squeezing is a standard technique in quantum optics and can be implemented under nowadays lab conditions. On the other hand, the cost of such an operation on infinite dimensional systems is not directly linked with the notion of gate complexity.
    Instead, we show that even skipping the amplification step (i.e., applying the identity gate to each mode), the whole protocol may still be faithful since the intensity gain is close to one. To see this, notice that for $g\geq 1$, ${}_{\rm coh}\langle g\cdot z|z\rangle_{\rm coh}=e^{-\frac{(g-1)^{2}}{2}|z|^{2}}$ for two coherent states. Therefore, the amplification error per mode scales as $O\left((g-1)^{2}\alpha_{0}^{2}\right)$ as long as $(g-1)^{2}\alpha_{0}^{2}\ll 1$, where $(\alpha_{0})^{2}$ (as in Protocol 2) upper bounds the intensity per mode. Substituting in the explicit configuration (80) (here $g^{2}=N/(N-N_{0})$), we get that the single-mode error scales as $O(n^{2}\cdot N^{-\Delta/3})$. Since there are $O(n)$ modes in total, the total error is $O(n^{3}\cdot N^{-\Delta/3})$. Since $N=\Theta(n^{32/3+\gamma})$, the error is vanishing whenever $\Delta>\frac{27}{32+3\gamma}$. This is achievable when $\gamma$ is large enough (i.e., when $N$ is large enough compared to $n$). We remark that there are also other promising and interesting approaches, e.g., finite-dimensional approximation via spin-coherent states, that are to be tested in future works. Overall, the amplification step can also be realized both time-efficiently and gate-efficiently.
In conclusion, we have shown that all subroutines of the compression protocol (Protocol 2) can be realized time-efficiently in principle. It can also be realized gate-efficiently, if the gate complexity of discarding qubits can be neglected (see the discussion on the truncation subroutine). We note that there is definitely room for further improving the algorithm’s efficiency, and the minimum complexity of shallow-circuit compression remains an interesting and important question to be explored in future works.